# mO Revanth[#1]

### v1

@@ Appendix A Some current datasets on AI-generated texts
We do analysis in many current dataset on AI-generated text and summarized them in Table 3. Many dataset contains fine-grained labels and multi-generator, but all of analyzed datasets are monolingual – only in English. This problem makes us to collect and formulate a new dataset for multilingual, multi-domain and multilingual AI-generated text dataset to adapt the approach to real-world scenarios.
@@
Human-initiated, AI-continued
@@
CyberHumanAI
Najjar et al. (2025)
English Human-written Wikipedia 500
@@ B.1 Statistics
Our FAIDSet includes 83,350 examples, which are separated into three subsets: train, validation, and test with the ratio shown in Table 4. The data set also comprises various sources of human-written texts, which are described in Table 5.
Subset Human AI Human-AI collab.
@@
Table 4: Number of examples per label in subsets in FAIDSet dataset.
@@ B.2 Diverse Prompt Strategies
To avoid biasing our generated corpora toward any single style or topic, we employ a broad set of prompt templates when synthesizing AI-generated and human-AI collaborative texts. By varying prompt structure, content domain, and complexity, we ensure that the resulting outputs cover a wide distribution of writing patterns, vocabulary, and rhetorical devices. This diversity helps our detector generalize more effectively to real-world inputs.
@@
Concretely, we sample two out of five prompts for AI-generated texts and several categories of human-AI collaborative texts in Table 6:
@@
  • AI-polished: Texts that were originally written by a human and then lightly refined by an AI system to improve grammar, clarity, or fluency without altering the core content or intent.
@@
  • AI-continued: Texts where a human wrote an initial portion (e.g., a sentence or paragraph), and an AI generated a continuation that attempts to follow the original style, tone, and intent.
@@
  • AI-paraphrased: Texts that were originally written by a human and then reworded by an AI system to express the same meaning using different phrasing, possibly altering sentence structure or word choice while preserving the original message.
@@
By mixing prompts across these categories, we generate a balanced corpus that mitigates over-fitting to any one prompt pattern and better reflects the diversity of real user queries.
Category Student Thesis Paper Abstract
AI-generated You are an IT student who are writing the graduation thesis. In a formal academic tone, write this section of a thesis on       , clearly outlining the structure, then write the passage concisely. The original text:       . In clear, structured prose, draft the section for a thesis titled       , cite some related works you mentioned in the passage and highlighting the contribution. The original text:       . You are a computer scientist who are very familiar with abstract writing for your works based on the title. Craft a concise word_count-word abstract for a paper titled       , summarizing the problem statement, methodology, key findings, and contributions. The original text:       . Compose a word_count-word abstract for the paper       , ensuring it includes motivation, approach, results, and implications for future research. The original text:       .
AI-polished You are an IT student who are refining your work to make it more complete. Refine the following section excerpt for grammar, clarity, and academic style while preserving its original meaning and terminology. The original text:       . You are an IT student who are refining your work to make it more complete. Enhance the academic tone, coherence, and logical flow of this thesis section without altering technical content. The original text:       . You are a computer scientist who are very familiar with abstract writing and refining the written abstract. Improve the coherence, precision, and formal tone of this draft abstract without introducing new content. The original text:       . Improve the clarity and conciseness of this abstract paragraph while maintaining all original findings and terminology. The original text:       .
AI-continued You are an IT student who are writing your graduation thesis. Continue to write the section from this thesis excerpt for approximately word_count words, maintaining formal academic structure and style. The original text:       . You are an IT student who are writing your graduation thesis. Extend the section by adding supporting detailed information for a thesis on       . The original text:       . You are a computer scientist who are very familiar with abstract writing. Add some concise concluding sentences to this partial abstract that highlights implications for future research. The original text:       . Continue the abstract by writing a closing statement that underscores the study’s contributions and potential applications. The original text:       .
AI-paraphrased Paraphrase the following thesis section in a clear academic tone, preserving citations and technical terms exactly. The original text:       . Reword this thesis excerpt to improve readability and maintain its scholarly voice, keeping all references unchanged. The original text:       . Rephrase this abstract in formal academic English, maintaining all original citations and technical accuracy. The original text:       . Paraphrase this abstract paragraph to enhance clarity and flow, ensuring all technical terms and citations remain intact. The original text:       .
@@
Table 6: Samples of diverse prompt templates used to generate FAIDSet
Appendix C Analysis on similarity across models and model families
@@
To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis using various perspectives: N-grams distribution, text length patterns, and semantic embedding visualization. These analyses allow us to examine similarities both within the same model family and across different model families, leading to a robust understanding of LLM "authorship" characteristics.
C.1 Text Distribution Between AI Families
@@
We first examined the distribution of text length, measured in both word and character counts across outputs from five AI models: Llama-3.3-70B-Instruct-Turbo Dubey et al. (2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Team et al. (2024). Using 2000 arXiv prompt seeds, each model generated a corresponding output, and their length distributions were plotted.
@@
C.2 Text Distribution Between AI Models Within the Same Family
We performed N-grams frequency analysis on responses generated by three models within the Gemini family: Gemini 2.0, Gemini 2.0 Flash, and Gemini 1.5 Flash using 500 texts from the arXiv abstract dataset. Figure 5 highlights overlapping high-frequency tokens and similar patterns in word usage and phrase structure among the three models.
Despite minor differences in architectural speed (e.g., Flash vs. regular) or release chronology, the N-grams distributions show minimal divergence. Frequently used tokens, such as domain-specific terms and transitional phrases, appeared with nearly identical frequencies. This suggests that these models share similar decoding strategies and training biases, likely due to shared pretraining corpora and optimization techniques resulting in highly consistent stylistic patterns.

### v2

@@ Appendix A Some Current Datasets on AI-generated Texts
We analyze many current datasets on AI-generated text and summarize them in Table 3. Many datasets contain fine-grained labels and multi-generators, but all of the analyzed datasets are monolingual – only in English. This problem prompts us to collect and formulate a new dataset for multilingual, multi-domain, and multilingual AI-generated text, enabling us to adapt the approach to real-world scenarios.
@@
Human-initiated, -continued
@@ B.1 Statistics
Our FAIDSet includes 83,350 examples, which are divided into three subsets: train, validation, and test, with the ratios shown in Table 4. The data set also comprises various sources of human-written texts, which are described in Table 5.
Subset Human AI human–LLM collab.
@@
Table 4: Number of examples per label in subsets in the FAIDSet dataset.
@@ B.2 Diverse Prompt Strategies
To avoid biasing our generated corpora toward any single style or topic, we employ a broad set of prompt templates when synthesizing LLM-generated and human–LLM collaborative texts. By varying prompt structures, content domain, and complexity, we ensure that the resulting outputs cover a wide distribution of writing patterns, vocabulary, and rhetorical devices. This diversity helps our detector generalize more effectively to real-world inputs.
@@
Concretely, we use five prompts for LLM-generated texts in Table 6 and several categories of human–LLM collaborative texts in Table 7, Table 8 and Table 9 consecutively:
@@
  • LLM-polished: Texts that a human initially wrote and then lightly refined by an LLM system to improve grammar, clarity, or fluency without altering the core content or intent.
@@
  • LLM-continued: Texts where a human wrote an initial portion (e.g., a sentence or paragraph), and an LLM generated a continuation that attempts to follow the original style, tone, and intent.
@@
  • LLM-paraphrased: Texts that were initially written by a human and then reworded by an LLM system to express the same meaning using different phrasing, possibly altering sentence structure or word choice while preserving the original message.
@@
By mixing prompts across these categories, we generate a balanced corpus that mitigates over‐fitting to any one prompt pattern and better reflects the diversity of real user queries.
Student Thesis Paper Abstract
You are a university student majoring in computer science. Please briefly summarize the main idea of the following paragraph. After that, rewrite the paragraph based on this content. Write naturally in an academic style. The rewritten paragraph should be approximately the same length in characters as the original. The original text is:    In clear, structured prose, draft the section for a thesis titled   , cite some related works you mentioned in the passage, and highlight the contribution. The original text is:   . Please begin by briefly summarizing the main idea of the paragraph to ensure full comprehension and retention of all essential content. Then, rewrite the paragraph in a formal academic style, consistent with a university-level thesis. The rewritten section should read naturally, be coherent within the context of an academic paper, and have approximately the same character length as the original. Assume the role of a senior software engineer. Your task is to process a paragraph from a computer science thesis using a two-step method. Firstly, you must summarize by deconstruction, which involves analyzing the original text and providing a structured summary by identifying its primary purpose, key input parameters, and main outcomes. Secondly, rewrite the paragraph based on the structured summary. The new version must be technically precise, unambiguous, and logically structured, making it easy for another engineer to understand. The original text is:   . You are a research scientist preparing a paper for a top-tier computer science conference. The original text is:   . For the following paragraph from a thesis draft, you have to summarize the core contribution, which can be earned by beginning with providing a concise, one-sentence summary that captures the main scientific contribution or key finding of the paragraph. Then, rewrite the paragraph for publication by using the summary as a guide, ensuring it is written in a formal, objective, and precise tone suitable for a peer-reviewed publication. Ensure the rewritten text is information-dense yet easy for a fellow researcher to follow. You are a university student majoring in computer science. You need to extract the main idea by writing a summary of the paragraph’s main idea. Rewrite the paragraph based on the content you have summarized. The rewritten text should be in a formal academic style, read naturally, be coherent, and have approximately the same character length as the original. The original paragraph is:    Assume the role of a researcher with experience in writing abstracts for scientific papers. Write a short paragraph of approximately 150-200 words based on the topic conveyed by the provided file name. Start directly with the topic, presenting it clearly, objectively, and in a concise academic style. Use correct spelling and grammar, and write in a scholarly tone. Topic name:    You are a computer scientist who is very familiar with abstract writing for your works, based on the title. Craft a concise word_count-word abstract for a paper titled   , summarizing the problem statement, methodology, key findings, and contributions. The original text:   . Compose a word_count-word abstract for the paper   , ensuring it includes motivation, approach, results, and implications for future research. Act like a senior researcher acting as a peer reviewer. Your task is to analyze and then rewrite the provided abstract to improve its structure and clarity. Deconstruct the abstract: First, examine the original text and break it down into four key components: What is the core problem being addressed?; What is the proposed solution or methodology?; What were the key results of the experiments?; What is the main contribution or impact of this work? After that, you must reconstruct the information from the components using only the data from your deconstructed points, and synthesize a new, cohesive abstract of approximately 150-200 words. Original abstract:   . You are a research scientist specializing in the sub-field suggested by the paper’s title. Your task is to generate a plausible abstract based only on the title. Based on the title, first generate a bulleted list of the likely components this paper would cover: the specific problem it probably addresses, the methodology it might propose, the kind of results one would expect, and its potential impact or contribution to the field. After that, weave these hypothesized points into a compelling and professional abstract of 150-200 words. Write it as if you were the author, confidently presenting your work. Paper title:    You are a technical writer for a prestigious AI research blog. Your goal is to rewrite a standard academic abstract to make it more impactful and highlight its core breakthrough for a broader technical audience. First, read the original abstract and identify the single most important takeaway or the core breakthrough of the paper and summarize this in one sentence. Second, rewrite the abstract to be approximately 150-200 words in length. Start with a strong opening sentence that directly states the problem or the breakthrough you identified. Then, briefly explain the methodology and results, always connecting them back to why they are important. The tone should be highly professional but more engaging than a typical arXiv abstract. Preserve all technical terms and citations accurately. Paper title:   . Original abstract:   .
@@
Table 6: List of diverse prompt templates used to generate FAIDSet – Label: LLM-generated
Student Thesis Paper Abstract
You are a university student majoring in computer science who has been assigned the task of polishing the paragraph below, which is excerpted from an undergraduate thesis. Improve the paragraph to make it clearer, more coherent, and more precise, while maintaining the original author’s academic tone and writing style. Do not rephrase any reference materials, figure labels, or citations preserve them exactly as they appear in the original paragraph. The original text:   . You are a meticulous academic editor with a specialization in computer science theses. Your task is to polish the following paragraph for conciseness and impact. Focus on eliminating redundant words and phrases, replacing weak verb constructions with stronger, more active verbs, and ensuring each sentence contributes directly and efficiently to the paragraph’s central point. The core technical meaning, all specific terminology, citations, and references to figures must be preserved precisely. The original text:   . Assume you are an IT student who is refining your work to make it more complete. Refine the following section excerpt for grammar, clarity, and academic style while preserving its original meaning and terminology. Provide only the polished version, without any introductory or explanatory text. The original text:   . Act as a PhD candidate reviewing a section of an undergraduate thesis. Your primary goal is to enhance the logical flow and argumentative coherence of the paragraph below. Revise the text to ensure that sentences connect seamlessly with clear transitions. The paragraph should build a coherent argument from the opening sentence to the conclusion. Do not introduce new information or alter the original technical content, terminology, or references. Your focus is solely on restructuring the existing information to create a stronger, more persuasive, and logical narrative. The original text:   . You are a fourth-year IT student who is refining your work to make it more complete. Enhance the academic tone, coherence, and logical flow of this thesis section without altering technical content. The original text:   . You are a researcher who has been assigned the task of polishing the paragraph below, which is excerpted from an undergraduate thesis. Improve the paragraph to make it clearer, more coherent, and more precise, while maintaining the original author’s academic tone and writing style. Do not rephrase any reference materials, figure labels, or citations—preserve them exactly as they appear in the original paragraph. The original text:  . You are a scientist who is very familiar with abstract writing and refining the written abstract. Improve the coherence, precision, and formal tone of this draft abstract without introducing new content. Provide only the polished version, without any introductory or explanatory text. The original text:  . Improve the clarity and conciseness of this abstract paragraph while maintaining all original findings and terminology. Provide only the polished version, without any introductory or explanatory text. The original text:   . You are a program chair for a leading academic conference, skilled at identifying impactful research. Your task is to polish the following arXiv abstract to maximize its impact and make its core contribution immediately apparent. Focus on sharpening the opening sentence to act as a compelling hook. Rephrase and reorder sentences as needed to clearly convey the main findings and highlight the significance of the work. All original terminology, data, and citations must be strictly preserved. Provide only the polished version as a single, continuous paragraph. The original text:   . You are a meticulous editor for a top-tier scientific journal, such as Nature or Science. Your objective is to polish the following arXiv abstract to enhance its technical precision and information density. Scrutinize every word to ensure it is the most accurate choice. Refine phrasing to eliminate any ambiguity and, where supported by the text, replace qualitative descriptions with more specific, quantitative statements. The goal is a text where every clause delivers critical information efficiently. Do not alter the scientific findings, technical terms, or citations. The original text:   .
Table 7: List of diverse prompt templates used to generate FAIDSet – Label: LLM-polished
Student Thesis Paper Abstract
You are a university student majoring in computer science who has been assigned the task of continuing the content of the paragraph below, which is excerpted from an undergraduate thesis. Please continue the text naturally, striving to mimic the tone and writing style of the given paragraph to avoid any inconsistency in expression, while ensuring clarity and coherence in an academic style. Do not rephrase the reference materials, figure labels, or citations-preserve them exactly as they appear in the original paragraph. The original text:   . You are an IT student who is writing your graduation thesis. Continue to write the section from file name    thesis excerpt for approximately word_count words, maintaining formal academic structure and style. Do not rephrase the reference materials, figure labels, or citations—preserve them exactly as they appear in the original paragraph. The original text:   . Act like an IT student who is writing your graduation thesis. Extend the section by adding supporting detailed information for a thesis on   . Do not rephrase the reference materials, figure labels, or citations—preserve them exactly as they appear in the original paragraph. The original text:   . You are a computer science researcher meticulously documenting your work. Your task is to continue the paragraph starting with the sentence below. The continuation must elaborate on the underlying mechanism, process, or rationale implied by the initial statement, effectively answering the how or why. The completed paragraph should be logically sound and consistent with the topic of a thesis titled   . Maintain a formal academic tone and provide only the continuation as a single, seamless paragraph. Initial sentence:   . You are a final-year IT student analyzing your research findings for your graduation thesis. Continue the paragraph that begins with the key statement below by providing an analytical extension. Your writing should focus on comparing the statement to existing work, contrasting it with alternative approaches, or discussing its broader implications within the context of the thesis titled   . Ensure the analysis is coherent and maintains a scholarly tone. Initial statement:   . You are a researcher who has been assigned the task of continuing the content of the paragraph below, which is excerpted from an undergraduate thesis. Please continue the text naturally, striving to mimic the tone and writing style of the given paragraph to avoid any inconsistency in expression, while ensuring clarity and coherence in an academic style. Do not rephrase the reference materials, figure labels, or citations-preserve them exactly as they appear in the original paragraph. The original text:    You are a scientist who is very familiar with abstract writing. Add some concise concluding sentences to this partial abstract that highlight implications for future research. Do not rephrase the reference materials, figure labels, or citations-preserve them exactly as they appear in the original paragraph. The original text:   . Continue the abstract by writing a closing statement that underscores the study’s contributions and potential applications. Do not rephrase the reference materials, figure labels, or citations-preserve them exactly as they appear in the original paragraph. The original text:   . You are a research scientist continuing the draft of a paper’s abstract. The provided text introduces the core problem or context. Your task is to continue the abstract by providing a concise description of the proposed methodology or approach. Detail the key techniques, model architecture, or experimental setup used to address the problem, ensuring the description is plausible for a paper titled   . The continuation must seamlessly connect to the initial text to form a single, coherent paragraph. Provide only the new text. Paper title:   , initial text:   . You are the lead author of a scientific paper summarizing your work. The text below already outlines the problem and methodology. Your task is to continue the abstract by presenting the key results and findings. Report on the primary outcomes, important performance metrics, or significant observations derived from your experiments. The results must be specific, quantitative where possible, and logically follow from the described method. The output must integrate smoothly with the initial text to form a single, cohesive paragraph. Initial text:   .
Table 8: List of diverse prompt templates used to generate FAIDSet – Label: LLM-continued
Student Thesis Paper Abstract
You are an academic writing tutor. Your task is to perform a deep paraphrase of the following thesis excerpt. The goal is to create a version with significant structural and lexical differences from the original, while rigorously preserving the precise meaning, nuance, and all technical information. Focus on altering sentence construction and rephrasing ideas in a completely fresh way. All specific technical terms, citations, and figure labels must remain unchanged. The tone must remain formal and scholarly. The original text:   . You are a Computer Science Student tasked with paraphrasing the following paragraph, which has been extracted from the thesis of an undergraduate student. Paraphrase the given thesis content while preserving its original meaning and context. Maintain clarity, coherence, and an academic tone. Do not paraphrase: References, figure labels, and citations should remain unchanged. The original text:   . Paraphrase the following thesis section in a clear academic tone, preserving citations and technical terms exactly. Do not paraphrase: References, figure labels, and citations should remain unchanged. The original text:   . You are a senior researcher mentoring a student on their thesis. Paraphrase the following paragraph with the primary goal of improving clarity and directness. Untangle convoluted sentences and rephrase the content using a more straightforward structure. The aim is to express the same technical information in a way that is easier for a reader to parse, without losing any nuance or academic rigor. Do not alter or rephrase technical terminology, citations, or references to figures. The original text:   . Reword this thesis excerpt to improve readability and maintain its scholarly voice, keeping all references unchanged. Do not paraphrase: References, figure labels, and citations should remain unchanged. The original text:   . You are a researcher with paraphrasing the following paragraph, which has been extracted from the abstract of a science paper in the computer science domain. Paraphrase the given abstract content while preserving its original meaning and context. Maintain clarity, coherence, and an academic tone. Do not paraphrase: References, figure labels, and citations should remain unchanged. The original text:    Rephrase this abstract in formal academic English, maintaining all original citations and technical accuracy. Do not paraphrase: References, figure labels, and citations should remain unchanged. The original text:   . You are an expert scientific editor tasked with reframing an abstract to maximize its immediate impact. Perform a structural paraphrase on the following text. First, identify the core components of the abstract (Problem, Method, Results, Contribution) internally. Then, rephrase and reorder these components to lead with the main Contribution or Result, followed by the problem it solves and the method used. This inverted structure should create a fresh and compelling narrative while preserving all original information. Strict requirements: All technical terms, data, and citations must be preserved exactly. The original text:   . Paraphrase this abstract paragraph to enhance clarity and flow, ensuring all technical terms and citations remain intact. Do not paraphrase: References, figure labels, and citations should remain unchanged. The original text:   . You are a senior scientist adapting a specialized abstract for a broader scientific audience. Your task is to paraphrase the following text to make it more accessible to researchers in adjacent fields, without sacrificing technical rigor. Rephrase the abstract by substituting hyper-specific jargon with more widely understood technical equivalents, but only if the precise meaning is retained. The goal is for a researcher from a different subfield to quickly grasp the core concepts. Strict requirements: The abstract’s original meaning, key findings, and data must be perfectly preserved. All citations and figure references must remain unchanged. The original text:
Table 9: List of diverse prompt templates used to generate FAIDSet – Label: LLM-paraphrased
Appendix C Experimental Details and Real-World Use Cases
C.1 Experimental Setup
Unless noted otherwise, we use a batch size of 64, the AdamW optimizer with a learning rate of 2e52e-5 and weight decay of 1e41e-4, 50 epochs, and 2000 warm-up steps. For the fuzzy kNN component, we set top-K = 20. We use temperature = 0.7.
C.2 Computational Cost
All experiments run on a single NVIDIA A100 (40 GB); total wall-clock training time for FAID on FAIDSet is approximately 5 hours.
In our setup using FAISS Douze et al. (2024) with CPU inference, the average query latency is approximately 5ms for FAIDSet with 84,000 embeddings on a standard server (Intel Xeon CPU, 64GB RAM). When performing with Fuzzy KNN, it takes some minor time to process, so the total average time for inference is 10ms.
C.3 Real-world Management of Vector Database
In the case of FAIDSet, our entire training and validation embedding store requires approximately 200MB, which is easily handled by standard server disks.
For larger datasets, we believe the system remains scalable:
  • FAISS supports disk-based indices and optimized search methods to keep memory usage low even with millions of vectors.
  • Modern servers with hundreds of GB of disk are sufficient for storing large embedding banks, and the use of CPU-only inference (since Fuzzy KNN is GPU-free) makes the architecture cost-effective for deployment in resource-constrained environments.
C.4 Generalizability of FAID on unseen human–LLM type of collaboration
To further assess FAID’s robustness to unseen collaborative writing patterns between humans and LLM systems, we conducted an additional experiment focusing on hybrid authorship styles that were not present in the original training dataset.
We performed a manual quality control process to enhance the reliability of our dataset and ensure the robustness of the FAID model. A team of five annotators, all IT-major university students fluent in English and aged 18-25, participated in the annotation process. Each annotator was assigned approximately 80 samples, covering all collaborative writing styles (LLM-generated, LLM-continued,LLM-paraphrased, and LLM-polished), resulting in a total of 400 human-reviewed instances representing various collaboration modes.
During the manual revision stage, annotators followed two key quality control principles:
  1. 1.
    Ensuring logical and informational consistency between the output samples and original texts, where the output length was controlled to be between 70%–150% of the source length.
  2. 2.
    Improving linguistic quality through spelling correction, synonym replacement, and contextual word refinement to ensure natural fluency and stylistic coherence.
FAID was evaluated on this manually curated dataset without any further fine-tuning. Despite these samples representing an unseen distribution, the model achieved a strong overall accuracy of 84.8%, precision of 82.8%, and recall of 85.0%, across all collaboration categories. This suggests a strong generalization capability beyond the data distributions encountered during model development. It also highlights FAID’s sensitivity to fine-grained stylistic blending, where human revision only partially conceals the generative footprint of LLMs.
C.5 Generalizability of FAID on real-world scenarios.
To further assess the generalizability of FAID beyond the controlled benchmark setting, we conducted an additional user study simulating realistic academic and professional writing scenarios. The goal was to evaluate whether FAID can maintain detection accuracy when applied to authentic, unconstrained text produced through real interactions with LLMs.
A group of five volunteers with diverse academic backgrounds (same as in the Appendix C.4). Each participant was instructed to engage with four popular AI systems: ChatGPT, Gemini, DeepSeek, and Llama 3.1 on Together.AI to generate text resembling authentic student or professional writing tasks. The interactions were open-ended but guided by three types of prompts commonly encountered in real-world use cases: writing a paper summary (as an abstract) and writing a passage for their own graduation thesis. Each participant was asked to generate five outputs per model, yielding a total of 200 real-world text samples (5 participants $\times$ 4 models $\times$ 10 trials). To ensure authenticity and diversity, participants were encouraged to adjust prompts iteratively, include follow-up clarifications, and perform light editing, mimicking realistic human–LLM co-writing behavior.
FAID achieved strong performance in this real-world setting, with an overall accuracy of 88.5%, precision of 85.9%, and recall of 89.7%, indicating that FAID still exhibits strong generalizability in uncontrolled real-world scenarios with in-domain settings.
Appendix D Analysis on Similarity across Models and Model Families
@@
To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis from various perspectives, including N-gram distribution, text length patterns, and semantic embedding visualization. These analyses allow us to examine similarities both within the same model family and across different model families, leading to a robust understanding of AI "authorship" characteristics.
D.1 Text Distribution between LLM Families
@@
We first examined the distribution of text length, measured in both word and character counts across outputs from five LLMs: Llama-3.3-70B-Instruct-Turbo Dubey et al. (2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Team et al. (2024). Using 2000 arXiv prompt seeds, each model generated a corresponding output, and their length distributions were plotted.
@@
D.2 Text Distribution Between LLMs Within the Same Family
We performed N-gram frequency analysis on responses generated by three models within the Gemini family: Gemini 2.0, Gemini 2.0 Flash, and Gemini 1.5 Flash using 500 texts from the arXiv abstract dataset. Figure 5 highlights overlapping high-frequency tokens and similar patterns in word usage and phrase structure among the three models.
Despite minor differences in architectural speed (e.g., Flash vs. regular) or release chronology, the N-gram distributions show minimal divergence. Frequently used tokens, such as domain-specific terms and transitional phrases, appeared with nearly identical frequencies. This suggests that these models share similar decoding strategies and training biases, likely due to shared pretraining corpora and optimization techniques resulting in highly consistent stylistic patterns.